The product of two numbers is 88. Algebraically determine the numbers if one number is 3 more than the other.
Answer by JulietG(1812)

8 and 11
The way I did it was finding the factors of 88 (A * B = 88) and then looking at which pair fit the equation A = B + 3
Well, sort of. Honestly, this is a 2+2=4 question...